OpenStudy (anonymous):

Find the derivative of y with respect to the appropriate varible. y=cos^-1(1/x)

OpenStudy (anonymous):

\[\cos^{-1} (1/x)\]

OpenStudy (anonymous):

Is the answer... \[1 / (1-(1/x)^2)\]

OpenStudy (noelgreco):

Remember \[\ \frac{ d }{ dx }\arccos x = \frac{ 1 }{ \sqrt{(1-x ^{2}} }\], and don't forget the chain rule.

OpenStudy (anonymous):

How do I use the chain rule for this problem?

OpenStudy (anonymous):

\[1/\sqrt{1-(1/x)^2(2(1/x)(-1)}\]

OpenStudy (anonymous):

Anyone?

OpenStudy (turingtest):

let u=1/x-->u'=? y=cos^-1(u)

OpenStudy (turingtest):

the chain rule gives\[D_xy=\frac1{\sqrt{1-u^2}}\cdot u'\]

OpenStudy (turingtest):

there is no "u" aside from the fact that I made it up to illustrate where the chain rule comes in

OpenStudy (anonymous):

Oh, i see. \[1/\sqrt{1-(1/x)^2} * -1\]

OpenStudy (turingtest):

\[\frac d{dx}(\frac1x)\neq-1\]

OpenStudy (turingtest):

\[\frac d{dx}(\frac1x)=\frac d{dx}(x^{-1})=?\]

OpenStudy (anonymous):

-1(x^-2)

OpenStudy (turingtest):

yes, that is what I called u'

OpenStudy (anonymous):

\[1/\sqrt{1-(1/x)^2} * -1(x^-2) \]

OpenStudy (anonymous):

\[-x^-2/\sqrt{1-(1/x)^2}\]

OpenStudy (turingtest):

yes, which can be written\[-\frac1{x^2\sqrt{1-\frac1{x^2}}}\]or a few other ways

OpenStudy (anonymous):

But either way is fine right?

OpenStudy (anonymous):

Thanks, I think I understand this now.